OpenStudy (anonymous):

I need help with Geometric Series 1+2+4+8 + ... a10

OpenStudy (michele_laino):

we can apply this formula: \[{a_{10}} = {a_1}{q^9}\] being \(a_1\) the first term, namely \(a_1=1\) and \(q= a_3/a_2=2\). \(q\) is the constant of your geometrical series So, substituting, we have: \[{a_{10}} = {a_1}{q^9} = 1 \cdot {2^9} = ...?\]

OpenStudy (anonymous):

=512